OpenStudy (anonymous):

Geometric Series PLEASE HELP FAST!!! MEDALL Ball 3 Height 1 3 ft (starting point) Height 2 0.9 Height 3 0.3 Height 4 0.1 Height 5 0.02 What is the average common ratio between the successive height values of ball 3?

OpenStudy (cwrw238):

you can easily work these out 0.9/3 , 0.3 / 0.9 etc

OpenStudy (anonymous):

im not good with problems like this.

OpenStudy (cwrw238):

then find the average of the 4 results

OpenStudy (cwrw238):

you can work out 0.9 / 3 right?

OpenStudy (anonymous):

sorry it took so long i had to make tea, but its 0.3.

OpenStudy (anonymous):

@Abhisar can you help me figure out how to do this? do i do it: 3/.09 0.9/0.3 0.3/0.1 0.1/0.02

OpenStudy (cwrw238):

right its 0.3 and 0.3 / 0.9 = 0.333 0.1 / 0.3 = 0.333 0.1 / 0.02 = ?

OpenStudy (cwrw238):

sorry the last one is 0.02 / 0.1

OpenStudy (cwrw238):

can you do that use your calculator if you like

OpenStudy (anonymous):

0.02

OpenStudy (cwrw238):

no that 0.02 divided by 1 its 0.02 divided by 0.1

OpenStudy (anonymous):

0.2.. oops

OpenStudy (cwrw238):

right so now you have 4 values and you need to find their average so add up these 4 values and divide by and you have your answer (0.3 + 0.333 + 0.333 + 0.2 ) / 4

OpenStudy (anonymous):

so is it 0.7415?

OpenStudy (cwrw238):

no what do you get the sum of the 4 numbers?

OpenStudy (anonymous):

1.166

OpenStudy (cwrw238):

right so divide this number by 4 to get the average

OpenStudy (anonymous):

is it 0.2915?

OpenStudy (cwrw238):

yes though 0.29 would be an ok figure

OpenStudy (cwrw238):

2 decimal places would be sufficient

OpenStudy (anonymous):

okkaaayy!! THANK YOUUU!!!

OpenStudy (cwrw238):

yw